OpenStudy (anonymous):

Help please i don't know what to do ;-; Two lines, A and B, are represented by the following equations: Line A: 2x + 2y = 8 Line B: x + y = 3 Which statement is true about the solution to the set of equations? It is (1, 2). There are infinitely many solutions. It is (2, 2). There is no solution.

OpenStudy (anonymous):

@nincompoop @pooja195

OpenStudy (anonymous):

@ganeshie8 @triciaal

OpenStudy (triciaal):

the solution is the set of values (x,y) that satisfies both equations at the same time did you sketch the graph and see where they meet?

OpenStudy (triciaal):

this is even easier divide the first equation by 2 to show x + y = 4 the second says x + y = 3 x + y = 4 cannot =3 at the same time so no solution